OpenStudy (anonymous):

A map is scaled so that 3 cm on the map is equal to 21 actual miles. If two cities on the map are 5 cm apart, proportion would you use to solve the problem? A. map/actual = 21/3 = 5/x B. map/actual = 3/21 = 5/x C. map/actual = x/21 = 5/3 D. map/actual = 3/5 = 21/x

OpenStudy (anonymous):

another ratio... :)

OpenStudy (anonymous):

what is the ratio though ? there's 3 numbers .

OpenStudy (phi):

each answer says map to actual can you identify which numbers are with the map?

OpenStudy (anonymous):

yes, you know the ratio of "map to actual" which is 3 cm to 21 miles, or (3/21)

OpenStudy (anonymous):

so if two cities on the map are 5 cm apart, how far are they ACTUALLY apart? You use the ratio of (map/actual) and the one piece of info you have... that the cities are 5 cm apart on the map... and you solve to find out how far the same cities are ACTUALLY apart.

OpenStudy (anonymous):

This is like your Sonic question... the mini-sign had side length 4, and the ratio of (actual/mini) was (15/2). So you multiplied 4 x (15/2) to get the ACTUAL side of 30. Here, you have a ratio, (3/21) or (21/3). You need to set it up correctly so that when you multiply 5 cm times the ratio, you get the right number of ACTUAL miles.

OpenStudy (anonymous):

\[\frac{ 5 cm }{ 1 } \times \frac{ 21miles }{ 3 cm } = \frac{ (5)(21) }{ 3 } = 35\]

OpenStudy (anonymous):

35 miles, that is...

OpenStudy (anonymous):

so option C ?

OpenStudy (phi):

they want map over actual 5/3 is not map over actual. it is map over map

OpenStudy (phi):

read this sentence 3 cm on the map is equal to 21 actual if you make it a ratio (just keep the numbers) you get part of the answer

OpenStudy (anonymous):

so A would be the best option

OpenStudy (anonymous):

no... but it is confusing, isn't it!! The way they are setting this up, all the ratios are "map / actual" (map / actual) = (3 / 21) = (5 cm / x miles)

OpenStudy (phi):

it has to be in the correct order!

OpenStudy (anonymous):

So it's option B... do you see why?